# Neutron Tagging Can Greatly Reduce Spallation Backgrounds in Super-Kamiokande

### v1

@@ Appendix C Comparisons to scintillator detectors
Our proposal to reduce spallation backgrounds by neutron tagging of showers is significantly different from a technique first developed by Borexino and used for other scintillator detectors [39, 41, 42]. This technique is designed to reduce backgrounds due to 11C, an important spallation product. This isotope can be produced by a gamma ray from a muon-induced electromagnetic shower initiating a $(\gamma,n)$ interaction. For possible 11C beta-decay events, one evaluates a triple coincidence between the possible decay event and the distances to the muon and the nearest neutron capture. In contrast, our proposal is based on the presence of neutrons due to hadronic showers, preceding isotope production, supported by the extensive theoretical work in this paper. And it is designed to cover all isotopes.

### v2

@@ Appendix C Comparisons to scintillator detectors
Our proposal to reduce spallation backgrounds by neutron tagging of showers is significantly different from a technique first developed by Borexino and used for other scintillator detectors [39, 40, 41, 42]. This technique is designed to reduce backgrounds due to 11C, an important spallation product. This isotope can be produced by a gamma ray from a muon-induced electromagnetic shower initiating a $(\gamma,n)$ interaction. For possible 11C beta-decay events, one evaluates a three-fold coincidence between the possible decay event and the distances to the muon and the nearest neutron capture. In contrast, our proposal is based on the presence of neutrons due to hadronic showers, preceding isotope production. Recently, Borexino used a neutron-multiplicity technique to improve on the three-fold coincidence veto for 11[41]. In related work, KamLAND-Zen utilized neutron multiplicities for spallation studies on xenon [42] Our approach is significantly more general than these — e.g., in that it applies to many isotopes — and is supported by the extensive theoretical work above.